### v1

@@ 4.2 Evidential epistemic uncertainty
We propose here to extend the notion of epistemic uncertainty to rich labels, by removing the dependence on observations, simplifying the computational phase, and allowing the model to detect new areas of uncertainty.
The epistemic uncertainty can be extended to rich labels by using the notion of plausibility within the framework of belief functions. It represents the total evidence that does not support the complementary event for a class $\omega$ or more generally for an element $A\in 2^{\Omega}$. The plausibility $Pl$ defines the belief that could be allocated to $A$:
@@
We have $Pl(A)=1-Bel(\bar{A})$. Analogous to equation (8) and for two classes $\Omega=\{0,1\}$ the epistemic uncertainty is maximal when both classes are highly plausible. The proposed evidential epistemic and aleatoric uncertainties are defined as follows:
@@
Figure 10: On the left, a sample of a dataset of three classes with an area of ignorance (labeled with imprecision) and on the right areas of uncertainty according to non-evidential uncertainty sampling.
(a) Non-specificity
(b) Discord
(c) Klir
Figure 11: Areas of uncertainty corresponding to the datasets 9(a) according to the non-specificity, the discord and the total Klir uncertainty.
(a) Green class
(b) Red class
(c) Blue class
Figure 12: Areas of uncertainty corresponding to the datasets 9(a) according to the evidential epistemic uncertainty for green, red and blue classes.
(a) Evidential aleatoric
(b) Evidential epistemic
(c) Evidential total
Figure 13: Areas of uncertainty for evidential epistemic and aleatoric uncertainties, according to 9(a).
@@
The example in Figure 10 shows a dataset of three classes with a zone of ignorance for some labels (between the green and red classes). Probabilistic (4)-(5) and epistemic (8) uncertainties cannot model the imprecision present in the labels, this less complete uncertainty zone is represented in 9(b).
@@
The previous uncertainty resulting from the sum of the discord and the non-specificity is presented in Figure 11. It manages both exploration 10(a) and exploitation 10(b) to give a better representation of the uncertainty 10(c).
@@
The extension of the epistemic uncertainty, also introduced in this paper, is presented in the following experiments. First, the evidential epistemic areas of uncertainties for each of the three classes are presented in Figure 12. Then, the resulting evidential epistemic uncertainty of the model is deducted from equation (19) in Figure 13 along with the evidential aleatoric and total uncertainties.
5 Sampling on real world dataset
@@
Some datasets have been labeled in an uncertain and imprecise way by users during crowdsourcing campaigns [22]. We therefore have access to really imperfectly labeled datasets with rich labels. Conventional methods for computing model uncertainty do not take into account the degrees of imprecision of these rich labels. The two proposed methods are illustrated on Credal Dog-2, one of these datasets. Figure 14 shows the dataset on the two first components of a Principal Component Analysis. This is a two-class dataset represented in 13(a) with true classes and in 13(b) with uncertain and imprecise rich labels given by contributors. Darker dots indicate higher certainty, and vice versa.
@@
Figure 16 shows the result of the first proposed method, sampling by Klir uncertainty, on the dataset with rich labels. The non-specificity is presented 14(a) and can be interpreted as the ignorance zones of the model. Discord is also represented 14(b) and the total uncertainty 14(c) is the sum of the two, it is this latter information that is used to sample on the model uncertainty.
@@
The second proposed method, the extension of epistemic uncertainty, which is a reducible uncertainty applied to evidential reasoning, is presented in Figure 16. The irreducible aleatoric evidential uncertainty 15(a) is presented along with the reducible epistemic evidential uncertainty 15(b). The total uncertainty 15(c) is the sum of the reducible and irreducible uncertainties. For active learning, it is not the total uncertainty, but the epistemic reducible uncertainty that is used.
6 Discussion & Conclusion
The calculation of epistemic uncertainty (non-evidential) is demanding, and not necessarily accessible. It is, depending on the observations, necessary to go through several phases of computation, estimation of likelihood, maximum likelihood and optimization.
In this paper, we have proposed two new uncertainty sampling strategies and a new way to represent them. With these two proposed methods, the use of Klir uncertainty and the extended evidential epistemic uncertainty, a simple calculation on the output of the model allows to obtain the uncertainties. The objective is to also take into account the uncertainty present in richer labels, which is currently not possible. The first strategy is based on Klir’s uncertainty, combining discord (how self-conflicting the information is) and non-specificity (how ignorant the information is) in the model output. The second strategy extends epistemic (reducible) uncertainty to the evidential framework and to several classes, simplifying the computational phase.
@@ 6 Discussion & Conclusion
Figure 14: Credal Dog-2 dataset, Brittany breed is in green and Beagle in red.
(a) Non-specificity
(b) Discord
(c) Klir
Figure 15: Areas of uncertainty corresponding to the dataset 13(b) according to the non-specificity, the discord and to the total Klir uncertainty.
(a) Evidential aleatoric
(b) Evidential epistemic
(c) Evidential total
Figure 16: Areas of evidential epistemic uncertainty corresponding to 13(b).
This simplicity obviously has a counterpart: the model must be able to deliver a mass function, to represent uncertainty and imprecision in the output. Such models exist but are not numerous, among them are the much quoted Evidential $K$-Nearest Neighbors [5], Evidential Decision Trees [6, 4], Evidential Random Forest and even Evidential Neural Networks [23]. The proposed methods are compatible with probabilistic models (since a probability is a special mass function) but the full depth of evidence modeling would be lost.
The novelty of this work lies in the representation of new information for uncertainty sampling, rather than in performance comparison. The next step is to apply these models to active learning, where the learning model has access to a very limited number of labeled observations, and must choose the most relevant observations to label in order to increase performance. The ability of the model to define these areas of uncertainty, and to categorize these uncertainties, is then relevant information.
Supplementary materials
Reproducibility
This section deals with the experiments presented in this document, with the different parameters for each experiment. Note that every datasets as well as the code in Python allowing to reproduce these experiments are available at: https://anonymous.4open.science/r/evidential-uncertainty-sampling-D453.
Uncertainty sampling
The three datasets
are generated according to the following separation line:
  1. 1.
    $y=-3x$
  2. 2.
    $y=\text{sin}(0.8x)$
  3. 3.
    $(x-0)^{2}+(y-0)^{2}=6$
The uncertainties
are computed on a grid of $60*50$ points (it is the same for every experiments). The used model is $K$-NN ($K$-Nearest Neighbors), with a probabilistic output and on the distance-weighted version available with scikit-learn [16] (every other parameters are the scikit-learn default parameters). The model is trained on the dataset and then try to predict the class of every data point of the grid by using 10 nearest neighbors. The uncertainty used is the least confidence measure given in equation (5) and the same parameters are used on Fisher’s Iris dataset.
6.0.1 Example on real world dataset:
An example on real data is presented in Figure 17, Fisher’s Iris dataset is presented on two variables. The three Iris classes are shown in 16(a) and the areas of model uncertainty are shown in 16(b), these are the regions where the model would be the most uncertain to predict the class of a new observation. This is summarized by the fact that if a new observation falls into a red zone, the model will have more difficulty predicting its class.
(a) Iris Dataset
(b) Uncertainty
Figure 17: Fisher’s Iris dataset on two descriptive variables, petal length and width, with model uncertainties. Setosa in green, Versicolor in red and Virginica in blue.
Epistemic and aleatoric uncertainties
The dataset
is generated according to the separation line of equation $y=0$ and the data points are generated, with more points on the left according to a binary logarithm $\text{log}_{2}$.
The uncertainties
are computed according to the same version of the $K$-Nearest Neighbors algorithm. The experimental process follows the one described in [15]. The formulas are deduced from equation (6) as follows:
$\begin{split}&\pi(1|x)=\underset{\theta\in\Theta}{\text{sup}}\;\text{min}[\pi_{\Theta}(\theta),p_{\theta}(1|x)-p_{\theta}(0|x)],\\ &\pi(0|x)=\underset{\theta\in\Theta}{\text{sup}}\;\text{min}[\pi_{\Theta}(\theta),p_{\theta}(0|x)-p_{\theta}(1|x)].\end{split}$ (20)
That can be rewritten with $f(a)=2a-1$:
$\begin{split}&\pi(1|x)=\underset{\theta\in\Theta}{\text{sup}}\;\text{min}[\pi_{\Theta}(\theta),f(p_{\theta}(1|x))],\\ &\pi(0|x)=\underset{\theta\in\Theta}{\text{sup}}\;\text{min}[\pi_{\Theta}(\theta),f(1-p_{\theta}(1|x))],\end{split}$ (21)
with:
$\pi_{\Theta}(\theta)=\frac{L(\theta)}{L(\hat{\theta})},$ (22)
For a Parzen window, as described in [15] and with $p$ and $n$ the number of positive and negative instances within the window, $L(\theta)$ is:
$L(\theta)=\binom{p+n}{p}\theta^{p}(1-\theta)^{n},\;\text{and}\;\;\hat{\theta}=\frac{p}{p+n}$ (23)
As we are not using a Parzen window but the distance-weighted $K$-NN, we take $p$ the sum of the inverse distance for positive elements among the $K$ neighbors, and $n$ reciprocally the same for the negative class. Yet we have:
$\begin{split}&\pi(1|x)=\underset{\theta\in[0,1]}{\text{sup}}\;\text{min}\left(\frac{\theta^{p}(1-\theta)^{n}}{\left(\frac{p}{p+n}\right)^{p}\left(\frac{n}{p+n}\right)^{n}},2\theta-1\right),\\ &\pi(0|x)=\underset{\theta\in[0,1]}{\text{sup}}\;\text{min}\left(\frac{\theta^{p}(1-\theta)^{n}}{\left(\frac{p}{p+n}\right)^{p}\left(\frac{n}{p+n}\right)^{n}},1-2\theta\right).\end{split}$ (24)
Brent’s method is applied to find a local minimum (here the maximum) in the interval $\theta\in[0,1]$. Note that in this paper, the experimental results are not there to compare a performance gap, but to support a new approach to sampling by uncertainty. And, in our case, we have deduced the following formula more appropriate, so it is this one that we use in the presented experiments:
$\begin{split}&\pi(1|x)=\underset{\theta\in[0,1]}{\text{sup}}\;\text{min}\left(\frac{\theta^{p}(1-\theta)^{n}}{\left(\frac{p}{p+n}\right)^{p}\left(\frac{n}{p+n}\right)^{n}},2\theta-1\right),\\ &\pi(0|x)=1-\underset{\theta\in[0,1]}{\text{sup}}\;\text{min}\left(\frac{\theta^{p}(1-\theta)^{n}}{\left(\frac{p}{p+n}\right)^{p}\left(\frac{n}{p+n}\right)^{n}},1-2\theta\right),\end{split}$ (25)
and for the uncertainties:
$\begin{split}&\mathcal{U}_{e}(x)=\text{min}[\pi(1|x),\pi(0|x)]-0.5,\\ &\mathcal{U}_{a}(x)=1-\text{max}[\pi(1|x),\pi(0|x)].\end{split}$ (26)
The computational complexity, the dependence on observations, the optimization phase, the use of likelihood and maximum likelihood make the model heavier. In the next two experiments and thus for the proposed models, a simple output-dependent calculation replaces all these steps.
Discord and non-specificity
The dataset
is generated according to the separation line of equation $y=0$ and the data points are added, with no observations in the subspace verifying $(x-2.8)^{2}+(y-0)^{2}<5$ and with imprecise (or more ignorant) observations in the subspace verifying $(x-4.5)^{2}+(y-0)^{2}<5$.
The evidential uncertainties
@@ 6.0.1 Example on real world dataset:
Evidential epistemic uncertainty
The dataset
is generated according to three gaussian scikit-learn blobs [16] with imprecise (or more ignorant) observations in the subspace verifying $(2.3-x)^{2}+(3-y)^{2}<2$.
The evidential epistemic uncertainty
is computed according to the same Evidential $K$-Nearest Neighbors with $K=6$ nearest neighbors, $\beta=2$, $\alpha=1$ and the formula (19) is used to compute both the proposed evidential epistemic and aleatoric uncertainties.
Non-regression for advanced methods
This section shows that for the models presented, whether epistemic uncertainty or the proposition to use Klir uncertainty and evidential epistemic uncertainty, none is reductive regarding the (4) and (5) classical measures. Effectively it offers similar areas of uncertainty as the one used is section 2 when epistemic, aleatoric and evidential uncertainties are indistinguishable. Figures 18, 19 and 20 show the areas of uncertainty on the datasets introduced in Figure 1. The epistemic uncertainty, Klir uncertainty and the evidential epistemic uncertainty offers, roughly, the same areas of uncertainty as the one used in Figure 1.
Figure 18: From left to right, the areas of uncertainty corresponding to the datasets 0(a)0(b) and 0(c) according to the epistemic uncertainty.
Figure 19: From left to right, the areas of uncertainty corresponding to the datasets 0(a)0(b) and 0(c) according to Klir uncertainty resulting from the discord and the non-specificity.
Figure 20: From left to right, the areas of uncertainty corresponding to the datasets 0(a)0(b) and 0(c) according to the evidential epistemic uncertainty.
AGg2+06hmGgXq+j3+/DsixYlgVN03a9Xu8jgCNCyIegIAgx13Vfd7vdu+FweG8YRkjXdWy329+dTgeSJD3ieZ7RNO0VAXAPwDEAO5VKndi2fWrb9jWl9Esul6PZbDY9Go1OZ7PZ9z/lyuD3OozU2wAAAABJRU5ErkJggg==" alt="[LOGO]">

### v2

@@ 4.2 Evidential epistemic uncertainty
We propose here to extend the notion of epistemic uncertainty (c.f. section 3) to rich labels (c.f. section 2.1), by removing the dependence on observations, simplifying the computational phase, and allowing the model to detect new areas of uncertainty.
The epistemic uncertainty can be extended to rich labels by using the notion of plausibility within the framework of belief functions (which differs here from the one presented in section 3). It represents the total evidence that does not support the complementary event for a class $\omega$ or more generally for an element $A\in 2^{\Omega}$. The plausibility $Pl$ defines the belief that could be allocated to $A$:
@@
We have $Pl(A)=1-Bel(\bar{A})$. For example, suppose the only reliable evidence is that a picture depicts an animal, and a Dog is an animal. In this scenario, it is entirely plausible that the picture is a Dog (plausibility is 1), yet there is no direct belief that the picture is a Dog (belief is 0). Analogous to equation (8) and for two classes $\Omega=\{0,1\}$ the epistemic uncertainty is maximal when both classes are highly plausible. The proposed evidential epistemic and aleatoric uncertainties are defined as follows:
@@
(c) Non-specificity
(d) Discord
(e) Klir
(f) Green class
(g) Red class
(h) Blue class
(i) Ev. aleatoric
(j) Ev. epistemic
(k) Ev. total
Figure 9: A three-class dataset: representing label imprecision and uncertainty zones.
@@
The example in Figure 9 shows a dataset of three classes with a zone of imprecision for some labels (between the green and red classes). Probabilistic (4)-(5) and epistemic (8) uncertainties cannot model the imprecision present in the labels, this less complete uncertainty zone is represented in Fig. 9(b).
@@
The previous uncertainty resulting from the sum of the discord and the non-specificity is also presented. It manages both exploration, Fig. 9(c), and exploitation, Fig. 9(d), to give a better representation of the uncertainty, Fig. 9(e).
@@
The extension of the epistemic uncertainty, also introduced in this paper, is presented in the same figure. First, the evidential epistemic areas of uncertainties for each of the three classes are presented in Fig. 9(f), Fig. 9(g) and Fig. 9(h). Then, the resulting evidential epistemic uncertainty of the model is deducted from equation (19) in Fig. 9(j) along with the evidential aleatoric and total uncertainties.
5 Experiments
In this section, we conduct two types of experiments. The first is more theoretical, applying the two proposed methods to a dataset with real rich labels. The second is more traditional in active learning, comparing one of the methods with uncertainty sampling on several real-world datasets. The exploration-exploitation dilemma is addressed in this second part.
This aim of the first exploratory and non-comparative experiment is to demonstrates how information is mapped by the model. The second experiment offers a more traditional metric-based approach in active learning, allowing for a tangible comparison of methods using classical metrics used in this domain [24].
5.1 Sampling on real world dataset
@@
In this section we use datasets for which we have access to truly imperfectly labeled data with rich labels. This part is exploratory in nature and does not endorse the superiority of any method. Moreover, conventional methods for computing model uncertainty do not take into account the degrees of imprecision of these rich labels and only have access to hard labels. This paper proposes two methods capable of addressing this gap. They are illustrated on Credal Dog-2, a dataset labeled uncertainly and imprecisely by users during crowdsourcing campaigns [17]. Figure 10 shows the dataset, on the two first components of a Principal Component Analysis to represent this 42-variable dataset in 2D. This is a two-class dataset represented in Fig. 10(a) with true classes and in Fig. 10(b) with uncertain and imprecise rich labels given by contributors. Darker dots indicate higher certainty, and vice versa.
@@
The first proposed method, sampling by Klir uncertainty, is demonstrated on the dataset with rich labels, Fig. 10(b). The non-specificity is presented in Fig. 10(d), and can be interpreted as the zones of imprecision of the model, either because it has not had enough access to information (lack of data) or because the users who labeled these instances are themselves ignorant. Discord is also represented in Fig. 10(c), these are the areas where the model’s prediction is conflicting, meaning it is closest to its decision boundary. The total uncertainty in Fig. 10(e) is the sum of the two, it is this latter information that is used to sample on the model uncertainty.
@@
The second proposed method extends epistemic uncertainty, which is a reducible uncertainty applied to evidential reasoning. The irreducible aleatoric evidential uncertainty in Fig. 10(f) is presented along with the reducible epistemic evidential uncertainty in Fig. 10(g). The total uncertainty in Fig. 10(h) is the sum of the reducible and irreducible uncertainties. For active learning, it is not the total uncertainty, but the epistemic reducible uncertainty that is used. Similarities can be noted between discord and aleatoric uncertainty and between non-specificity and epistemic uncertainty. Additionally, the areas of total uncertainty are also similar.
One advantage of the methods proposed in this paper is their ability to account for the uncertainty already present in the labels (i.e. the uncertainty of the oracles). During labeling by human [^7] This representation also applies to labeling performed by a machine. (as in the Cedal Dog-2 dataset), a user may hesitate between class 1 and 2. In such cases, it is preferable to model their uncertainty rather than forcing them to provide a wrong label for one of the classes, which would introduce noise into the dataset. The advantage of belief functions is that for multiple classes, the user can respond with various degrees of ignorance. For instance, they may indicate that they are uncertain about the true class but are relatively confident that it is not class 4 or 2. Therefore, being able to represent the imperfections of the oracle can also lead to improved results in machine learning. Unfortunately, extremely expressive visual results are hard to obtain (and interpret) since these datasets are rare, difficult to collect, and very noisy. However, the following experiment, in section 5.2, demonstrates that the proposed methods perform very significantly on this dataset.
@@
(c) Discord
(d) Non-specificity
(e) Klir
(f) Ev. aleatoric
(g) Ev. epistemic
(h) Ev. total
@@
Figure 10: Ignorance mapping in the Credal Dog-2 dataset: Brittany breed (green) and Beagle breed (red).
5.2 Application to active learning
Sampling by Klir uncertainty was chosen for this series of experiments, and the only parameter $\lambda$ in the proposed method is at the heart of this study. A $\lambda$ that tends towards 00 implies more exploitation, whereas a $\lambda$ that tends towards $1$ implies more exploration. For these experiments, we set $\lambda$ to 0.2, which means that the model will opt for more exploitation than exploration (this choice is motivated in section 6). Below, the results are compared with random sampling (the baseline) and the very popular uncertainty sampling (5).
The experiments [^8] Experiments where conducted according to the following code: https://anonymous.4open.science/r/evidential-active-learning-B266 have been carried out on datasets containing between 2 and 8 classes with a number of observations in different ranges. We used well-known datasets available on the UCI Machine Learning Repository [11], and very often used in active learning, as well as Dog-2, the dataset presented in the previous section. Table 1 describes these datasets, with the number of observations (or instances), the number of classes, the number of features, and the entropy for the distribution of classes [^9] An entropy of $1$ means that the classes are perfectly equidistributed and an entropy of 00 would indicate the total over-representation of one of the classes..
Table 1: Datasets description, with number of observations, classes, features and class distribution entropy.
Dataset Observations Classes Features Entropy
Bank 1372 2 4 0.99
Qsar 1055 2 41 0.92
Blod 748 2 4 0.79
Breast Cancer 569 2 30 0.95
Ionosphere 351 2 34 0.94
Heart 303 2 7 1.00
Liver 345 2 6 0.98
Sonar 208 2 60 1.00
Parkinson 195 2 22 0.81
Dog-2 200 2 42 1.00
Seeds 210 3 7 1.00
Iris 150 3 4 1.00
Wine 178 3 13 0.99
Glass 214 6 9 0.83
Ecoli 336 8 7 0.73
@@ 5.2 Application to active learning
Since the goal of active learning is to reduce the cost of labeling, one experiment involves evaluating the model’s performance as observations are progressively labeled. Experiments are arbitrarily stopped once the dataset has been labeled at $60\%$ (it will be clear from the graphs that there is no point in going any further). The model is the same for each method, the Evidential $K$-Nearest Neighbors of [10], with $K=7$ neighbours (see [15] for parameter selection). Each experiment is performed 100 times to obtain an estimation of the actual mean accuracy of the model for each dataset. Several criteria are used to compare the results, including accuracy, the area under the accuracy curve (AUAC) and the rank obtained for each dataset. For evaluation, several statistical tests are conducted, including Student’s t-test for AUACs, Friedman’s test, and the Wilcoxon-Holm method for critical difference diagrams [7].
(a) Dog-2
(b) Ionosphere
(c) Parkinson
(d) Ecoli
(e) Sonar
(f) Heart
Figure 11: Mean accuracy vs. number of labeled instances for Random Sampling, Uncertainty Sampling, and the proposed method with $\lambda=0.2$ for 6 datasets.
Figure 11 shows 6 of the 15 datasets where the proposed method offers the most significant performance, the final performance on the full labeled dataset (i.e. if there is no active learning) is represented by the dotted curve and the dashed curve represents the proposed method with $\lambda=0.2$. In each graph, the superiority of the proposed method over uncertainty sampling is evident, particularly in terms of AUAC. Notably, for the Sonar and Heart datasets, this superiority is only temporary, observed mainly at the beginning of active learning for Sonar and in the middle for Heart. Nevertheless, it will be demonstrated that this domination is not always statistically confirmed, especially for the Parkinson dataset. Assuming an identical labelling cost for each observation, some insights can be reported below.
Some active learning insights on Figure 11
  • On Dog-2: When reaching $99\%$ of the full dataset performance, uncertainty sampling manages to reduce labeling cost by $62\%$ whereas the proposed method manages to reduce the cost by $82\%$.
  • On Ionosphere: Using the proposed method, the labeling costs can be reduced by a factor of $9$ with $0\%$ of accuracy loss with respect to the full dataset, whereas with uncertainty sampling, to allow labeling $9$ times cheaper, the model would loose $6\%$ of accuracy (for this dataset, the reduction in labeling cost can improve the performance of the model, a phenomenon which sometimes occurs in active learning, represented by the active learning curve which exceeds the full dataset performance horizontal line).
  • On Ecoli: It takes $10$ steps for uncertainty sampling to reach the performance of the proposed method after only $3$ steps.
Even on datasets where the performance of the proposed method is lower, the gap is not always wide. Figure 12 shows two such datasets where the method performs less well. The difference is slightly greater between the proposed method and uncertainty sampling on the Iris dataset.
(a) Iris
(b) Wine
Figure 12: Mean accuracy vs. number of labeled instances for Random Sampling, Uncertainty Sampling, and the proposed method with $\lambda=0.2$ for Iris and Wine.
Since the objective here is to minimize labeling costs [^10] Although it can also be to maximize performance given a cost, one can set a performance threshold and focus on the actual reduction in costs. By accepting a 2% loss in performance (threshold of 98% performance on the full dataset), conventional uncertainty sampling reduces the number of labels in the Dog-2, Ionosphere, and Heart datasets by 76%, 86%, and 43% respectively. In contrast, the proposed method reduces the number of labels in the respective datasets by 88%, 91%, and 83%. However, a performance threshold has been set, which is why the area under the curve is a good indicator, as it captures the reduction in costs for all possible thresholds. Table 2 shows the mean areas under the curve for the three methods studied and for each dataset. A statistical t-test is also performed between the first and second best methods for each values. Random sampling performs best on the Liver dataset, uncertainty sampling performs best on Seeds, Iris and Wine and the proposed method performs best on the other 10 datasets, except Blod where there is a tie.
Table 2: Mean AUAC for random sampling, uncertainty sampling and the proposed method with $\lambda=0.2$ on each dataset. Student’s t-test is also performed to determine the significance of the best method.
Method t-test
Dataset Random Uncertainty $\lambda=0.2$ statistic p-value
Bank 81.17 81.49 82.23 2.74 0.0067
Qsar 97.69 99.15 99.16 0.35 0.7271
Blod 76.09 76.85 76.85 0.00 0.9965
Breast Cancer 93.87 94.96 95.31 1.84 0.0669
Ionosphere 75.77 81.06 82.40 2.33 0.0210
Heart 67.89 67.72 68.08 0.29 0.7741
Liver 58.07 57.37 58.02 0.07 0.9415
Sonar 67.79 70.67 70.94 0.37 0.7089
Parkinson 80.60 83.93 84.85 1.64 0.1034
Dog-2 90.94 93.06 94.10 3.24 0.0014
Seeds 88.70 89.93 89.48 0.84 0.4010
Iris 88.22 91.23 90.60 1.19 0.2373
Wine 91.66 93.55 93.27 0.86 0.3920
Glass 57.33 58.32 59.05 0.87 0.3829
Ecoli 78.59 80.89 81.98 2.19 0.0300
[…]
To statistically find the best model, a critical difference diagram is drawn up. The first diagram in Figure 13(a) is a comparison of the proposed method with different values of $\lambda$. On all datasets, $\lambda=0.2$ ranks on average at position $2.13$ out of $4$ and $\lambda=0.5$ (which is equivalent to as much exploration as exploitation) ranks on average at position $3.47$. If a line connects two methods, this means that despite the better performance of one, the methods are not statistically differentiable. In the example, $\lambda=0.2$, $\lambda=0.3$ and $\lambda=0.4$ are not statistically different. Now, Figure 13(b) is obtained by comparing the proposed method with random sampling and sampling by uncertainty. In average, the proposed method ranks $1.33$ out of $3$ and the significance of this result is demonstrated by the absence of a line linking the methods. It may also be interesting to note that uncertainty sampling and random sampling are not connected by a line either.
Figure 13: Critical difference diagrams for different values of $\lambda$ on the proposed method 13(a) and for random sampling, uncertainty sampling and the proposed method 13(b).
6 Discussion
Calculating epistemic uncertainty (non-evidential) is demanding and not always accessible. It depends on the observations, requiring several phases of computation, including likelihood estimation, maximum likelihood, and optimization.
The two proposed methods offer simplicity, but there is a counterpart: the model must be capable of delivering a mass function to represent uncertainty and imprecision in the output. Such models exist but are not abundant. Among them are Evidential $K$-Nearest Neighbors [10], Evidential Decision Trees [13, 9], Evidential Random Forests [16], and even Evidential Neural Networks [36]. The proposed methods are compatible with probabilistic models (since a probability is a special belief function), but they may not capture the full depth of evidence modeling.
In the experiments above, $\lambda$ was set at $0.2$, meaning that the model prioritizes exploitation over exploration. This is a value that gives good results. Our studies to determine when one is more relevant than the other are illustrated in Figure 13(a). The results indicate that several lambda values yield fairly similar outcomes. The value of 0.2 is the one that gives the best performance in general: for the majority of datasets, it is more interesting to do more exploitation, without exceeding a certain limit, otherwise model performance will drop. For future work, it would be interesting to modify the value of $\lambda$ as the labeling process progresses. This adjustment could lead to a more powerful model capable of dynamically balancing between exploration and exploitation.
7 Conclusion
This paper introduces two new uncertainty sampling strategies and a novel representation method for them. These two methods use Klir uncertainty and an extended evidential epistemic uncertainty. A straightforward calculation on the model output enables the extraction of uncertainties. The objective is to also take into account the uncertainty present in richer labels, which was not possible up to now. The first strategy is based on Klir’s uncertainty, combining discord (how self-conflicting the information is) and non-specificity (how imprecise the information is) in the model output. The second strategy extends epistemic (reducible) uncertainty to the evidential framework and to several classes, simplifying the computational phase.
The proposed Klir uncertainty sampling is chosen for its competitiveness in active learning. Its superiority over uncertainty sampling is statistically validated across the datasets examined in the experiments. The novelty of this work lies in representing new information for uncertainty sampling, which also yields significant performance improvements in traditional active learning. The next step is to control exploration and exploitation (represented here as the $\lambda$ parameter) and to determine, for each dataset, when exploration or exploitation is more advantageous. The ability of the model to define areas of uncertainty, and to categorize these uncertainties, is then relevant information.
Declarations
Funding
This work is funded by the Brittany region and the Côtes-d’Armor department. The authors also received funding from IRISA, the University of Rennes, DRUID and Orange SA.
Competing interests
Arthur Hoarau, Jean-Christophe Dubois, Yolande Le Gall and Arnaud Martin received research support from the University of Rennes, the IRISA laboratory and the DRUID team. Vincent Lemaire received research support from Orange SA.
Ethics approval
Not Applicable
Consent to participate
All authors have read and approved the final manuscript.
Consent for publication
All authors approved the publication.
Availability of data and materials
All data used is this study are available publicly online. The datasets were extracted directly in the repositories available with the links in the folloing section.
Code availability
The code for theoretical experiments is available at this link: https://anonymous.4open.science/r/evidential-uncertainty-sampling-D453. Link to the code for the experimental part in active learning: https://anonymous.4open.science/r/evidential-active-learning-B266.
Authors’ contributions
Arthur Hoarau, Vincent Lemaire, Jean-Christophe Dubois, Yolande Le Gall and Arnaud Martin contributed to the manuscript equally.